OpenStudy (anonymous):

Check please? A line passes through (–1, 7) and (2, 10). Which is the equation of the line? (Points : 1) –3x + y = 4 –3x + y = 16 –x + y = 8 <<<<< –x + y = 12

OpenStudy (anonymous):

by eq y-y1=m(x-x1)

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

x1=-1 and x2=2

OpenStudy (anonymous):

y1=7 and y2 =10

OpenStudy (anonymous):

Was I right or wrong?

OpenStudy (anonymous):

@danish071996

OpenStudy (anonymous):

u were right

OpenStudy (anonymous):

yes you did it correct! :)

OpenStudy (anonymous):

@CrazyCountryGirl

OpenStudy (anonymous):

Thanks!